OpenStudy (anonymous):

how do you rotate a function to a certain angle? say y = x^2+1 to 45degree

OpenStudy (anonymous):

you can draw a new XY coodinate system and center the graph about the line y=x

OpenStudy (anonymous):

I think you want the new equation also huh?

OpenStudy (anonymous):

Well, that would require a bit more work...

OpenStudy (anonymous):

well yes, of course, so any help?

OpenStudy (kinggeorge):

One of a couple problems that I'm seeing, is that \(x^2+1\) isn't a function anymore if you rotate it even the tiniest bit.

OpenStudy (anonymous):

ok yes it would be broken into various funtions, so how are they?

OpenStudy (kinggeorge):

I suppose one option would be to simply change the basis/variables. You can create new variables s, t such that the s axis is 45 degrees off the positive x-axis and the t axis is 45 degrees off the positive y-axis. Then write you equation as \(t=s^2+1\) and convert back to x, y after that.

OpenStudy (kinggeorge):

You would still run into problems about the domain when converting though.

OpenStudy (kinggeorge):

I think your change of variables would look something like \[s=\sqrt{2x^2}\]\[t=\sqrt{2y^2}\]

OpenStudy (anonymous):

I don't think that will do

OpenStudy (kinggeorge):

Not that I disagree with you, but why won't it work?

OpenStudy (anonymous):

never mind, I have found an answer

OpenStudy (kinggeorge):

Out of curiosity, what was it?

OpenStudy (anonymous):

first there is: x^2 - 2xy + y^2 - x * sqrt(2) - y * sqrt(2) = 0 and when solved for y there is: y = [(sqrt(2) + 2x) +/- sqrt(4x^2 + 2 + 4 sqrt(2) x - 4x^2 + 4 sqrt(2) x)] / 2 or y = [sqrt(2) + 2x +/- sqrt(8 sqrt (2) x + 2)]/2.